40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 24 01:21:11 2011, kiwamu, Update, SUS, damping test 
    Reply  Mon Sep 26 14:21:07 2011, kiwamu, Update, SUS, ITMX ULSEN shows jumps ITMXspectra.pngITMX-sensors.png
       Reply  Mon Sep 26 15:54:46 2011, kiwamu, Update, SUS, ITMX ULSEN shows jumps 
          Reply  Mon Sep 26 16:42:08 2011, kiwamu, Update, SUS, ITMX ULSEN : fixed connection.pngITMXsensors.png
             Reply  Wed Sep 28 07:46:42 2011, steve, Update, SUS, Dsub connections at the back of 1X5 are fixed 
Message ID: 5546     Entry time: Mon Sep 26 15:54:46 2011     In reply to: 5544     Reply to this: 5547
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: ITMX ULSEN shows jumps 

Currently the damping of the ITMX suspension is intentionally disabled for the noise investigation.

Quote from #5544

 

However the spectra from ITMX showed a noise floor at very high level. This is because of strange jumps in the signal of the UL shadow sensor.

I will check some analog circuits for the UL sensor.

 

ELOG V3.1.3-